OpenStudy (anonymous):

Let \[ f(n)=\left\lfloor \sqrt{n}\right\rfloor +n \] Show that for every integer m greater or equal to 1, the sequence \[ f^{k}(m), \, k=0,1,2,3,\cdots \] has at least one square.

OpenStudy (anonymous):

k=1,2,3 f^2(x)= f(f(x) f^3(x)=f(f(f(x)))

OpenStudy (anonymous):

yes you have to show it for every n

OpenStudy (anonymous):

No, @mahmit2012 , you're showing it for all \(m\), there exists a \(k\) such that \(f^k(m)\) is a square.

OpenStudy (anonymous):

No the sequence f^k(m), k=1,2,3,4,... has a square for every m

OpenStudy (anonymous):

Is that not saying the same thing as what I just said? \[\forall n,\ \exists k \in\mathbb N,\ f^k(n) \text{ is square}\] Also, is \(n\) restricted to particular values, such as \(\mathbb N\) or \(\mathbb Z\)?

OpenStudy (anonymous):

@mahmit2012 In order to show that, you need to show that there exists some \(k\) such that \(f^k(n)\) is a square. If there didn't exist some \(k\), then there would not be at least one square in the sequence, which transcends all possible values of \(k \in \mathbb N\).

OpenStudy (anonymous):

The sequence\[ (f^k(m))_{k \ge1} \] has a square for every \[m \ge 1\]

OpenStudy (anonymous):

\[ f^8(7)= 36=4^2 \\ f^{10}(7+1)= 50\ne (4+1)^2 =25 \]

OpenStudy (blockcolder):

Ah, Putnam problems. They look easy, but they're deviously hard.

OpenStudy (blockcolder):

In this book where I've seen the problem, the author presented a conjecture that was not quite right: If \(m=n^2+b \text{where } 0 \leq b <2n+1\), then \(f^2(m)=k^2+(b-1)\). But he said that a similar conjecture is correct, and leads to the full solution. So the key is finding the correct conjecture.

OpenStudy (experimentx):

*

OpenStudy (anonymous):

if i suppose there is no m,k wich satisfies it and prove there is a solution,then for absurd i will prove it

OpenStudy (anonymous):

@eliassaab what you think?

OpenStudy (anonymous):

Sorry, k=0 is included and \[ f^0(m)=m \]

OpenStudy (anonymous):

It is enough to prove it for all integers between two given squares.Fix n. It is enough to prove the statement for all m such that \[ n^2 \le m \le (n+1)^2 \]

OpenStudy (anonymous):

It is enough to prove it for all integers between two consecutive squares

OpenStudy (anonymous):

You did not attend my solution that was right. Check your solution again that is wrong and not generalized solution.

OpenStudy (blockcolder):

It's not true for r=3: \[m=7=2^2+3 \text{ but } f^6(m)\neq (2+3)^2\]

OpenStudy (anonymous):

Suppose that \[ m=n^2 + r \] Let us distinguish two cases: case I: \[ 1\le r \le n\\ f(m)=n^2 +r + n \\ f^2(m)=n^2 + r + n +n, \text { since } n^2 + r + n < n^2 + 2n +1\\ f^2(m)=n^2 + 2n +1 -1 +r=(n+1)^2 + r-1 \] If r =1, then \[ f^2(m)\] is a square. In the general case \[f^{2 r}(m)\] is a square. case II: \[ n+1\le r \le 2n +1\\ f(m)=n^2 +r + n= \\ n^2 + r + n+n -n +1 -1=\\ n^2 + 2 n +1 + r-n -1=\\ (n+1)^2 + r-n-1 \] If r =n+1, then \[ f(m)\] is a square. If r> n+1, you can show that if k= 2(r-n-1)+1, then \[ f^k(m) \] is a square.